OpenStudy (anonymous):

Two angles are supplementary. The first angle as 20deg. more than the second angle. How many degrees are in each angle? (Work shown would be greatly appreciated)

OpenStudy (anonymous):

*has

OpenStudy (anonymous):

Supplementary means 180..

OpenStudy (anonymous):

140?

OpenStudy (anonymous):

First angle = 20 + Second angle: 20 + Second Angle + Second Angle = 180 2S = 160 Divide both the sides by 2 and find the second angle..

OpenStudy (anonymous):

I don't get it. :(

OpenStudy (anonymous):

200

OpenStudy (anonymous):

Wait, 200 + 180?

OpenStudy (anonymous):

See, let the first angle be : x And other angle be y: So, first angle = 20 more than Second Angle, x = 20 + y Now you know, x + y = 180 replace x by (20 + y) (20 + y) + y = 180 2y = 160 Now find y from here..

OpenStudy (anonymous):

I still don't get it.

OpenStudy (anonymous):

Where you are facing problem??

OpenStudy (anonymous):

Just the whole problem...

OpenStudy (anonymous):

Let us start from the start.. Let first angle be x and second angle be y.. Getting till here??

OpenStudy (anonymous):

Ok,I'm with ya.

OpenStudy (anonymous):

Now the relation given in question is: First Angle = 20 more than Second Angle Right??

OpenStudy (anonymous):

And the 2nd angle is 180?

OpenStudy (anonymous):

No.... The sum of the two angles first and second is 180...

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

Is it right?? First Angle(x) = 20 more than Second Angle(y) x = 20 + y (more than means +) It is given in question.. Read the question once again... Got till here??

OpenStudy (anonymous):

What is y?

OpenStudy (anonymous):

y we assume as second angle.. There are two angles, no?? first is x and second is y (assuming this)..

OpenStudy (anonymous):

Yes there's two angles. I have t find out how many degrees are in each angle.

OpenStudy (anonymous):

*to

OpenStudy (anonymous):

See, have you got the stuff I have given above?? That is:' x = 20 + y Till here you got or not??

OpenStudy (anonymous):

Ok,I understand.

OpenStudy (anonymous):

is y 160,and x is 180?

OpenStudy (anonymous):

Now very first statement of your question says Two angles are Supplementary.. It means: First Angle + Second Angle = 180 Till here you got??? Are you in a hurry to solve this?? Sorry your answer is incorrect...

OpenStudy (anonymous):

So just divided 180 by two? Not really(:

OpenStudy (anonymous):

Do you want the solution or just the answer???

OpenStudy (anonymous):

The solution,please.

OpenStudy (anonymous):

See, the sum of two angle is 180 then how could one angle be 180?? It means other angle is zero????? Don't be fast.. Go slowly.. See,, First Angle + Second Angle = 180 (given) It means : x + y = 180 Got till here???

OpenStudy (anonymous):

So you're saying that x is 180 & y is 0?

OpenStudy (anonymous):

I am explaining to you what you said earlier... Sorry dear you are not responding to my questions clearly instead you are creating questions and confusion of your own..

OpenStudy (anonymous):

I'm so confused. D;

OpenStudy (anonymous):

Did not you say x = 180 and y = 160 tell me on summing them do you get 180???

OpenStudy (anonymous):

I thought that was what it was. If it was 180 & 160, the sum would've been 340.

OpenStudy (anonymous):

See, there are two variables x and y which only can be solved if we have two equations..

OpenStudy (anonymous):

Ok,I get you.

OpenStudy (anonymous):

What do you mean by this statement "Two angles are supplementary" tell me..

OpenStudy (anonymous):

That's just have the question says.

OpenStudy (anonymous):

Yes but I am asking for the meaning of this line???

OpenStudy (anonymous):

I think it means that they're the same?

OpenStudy (anonymous):

Oh My God.. remember these two lines: 1. Complementary means 90 degrees.. 2. Supplementary means 180 degrees.. Do you remember this for lifetime??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Two angles are supplementary does not mean that they are both of 180... It means: SUM OF THE TWO ANGLES = 180 Getting or not??

OpenStudy (anonymous):

It means that together they're 180?

OpenStudy (anonymous):

Yes they together are of 180.. Now you are getting my point.. For example they can be: 90 + 90 160 + 20 140 + 40 50 + 130 But they together form 180 Getting now???

OpenStudy (anonymous):

Yes! :D

OpenStudy (anonymous):

Now it will make your first equation: Given : First Angle + Second Angle = 180 Let first angle be x and Second angle be y So, x + y = 180 ----------------------1 Till here you got or not??

OpenStudy (anonymous):

I was following you until the "x+y=180 ---------1" part.

OpenStudy (anonymous):

I asked did you get this or not??

OpenStudy (anonymous):

Part of it.

OpenStudy (anonymous):

what that means??

OpenStudy (anonymous):

Maybe I could just get the answer?

OpenStudy (anonymous):

do you want answer only??

OpenStudy (anonymous):

Yeah,I think it'd be better to move onto another question..

OpenStudy (anonymous):

Well,it'd be two answers.

OpenStudy (anonymous):

Sorry I can't give you answer.. By this you will leave this concept and you are also not able to do problems like these... Going to sleep.. Good Night dear...

OpenStudy (anonymous):

Thanks anyways night.

OpenStudy (anonymous):

I just don't wanna fail the test,yenno.

OpenStudy (anonymous):

Call @jim_thompson5910 may be he can explain it better than me...

OpenStudy (anonymous):

Mhkay,thanks{:

OpenStudy (anonymous):

It is just not about the test It is about learning about knowledge...

jimthompson5910 (jim_thompson5910):

Can you find me two numbers that satisfy BOTH of the following conditions A) They add to 180 AND B) One number is 20 more than the other ???

jimthompson5910 (jim_thompson5910):

Clearly 180 and 0 add to 180....BUT...180 is NOT 20 more than 0 So 180 and 0 are NOT the answers

OpenStudy (anonymous):

I doubt I'm ever going to have to use this later in life,so whatever. No. ._.

OpenStudy (anonymous):

May God Bless You @Sabreena7898

OpenStudy (anonymous):

100 & 80?

jimthompson5910 (jim_thompson5910):

You got it

jimthompson5910 (jim_thompson5910):

You may use this later in life, you can never say "never"

jimthompson5910 (jim_thompson5910):

Do you know how to get the answer in general?

OpenStudy (anonymous):

I really don't see where this would come into play,but yes I do.Thanks(:

jimthompson5910 (jim_thompson5910):

alright great, and it's hard to see what we'll need down the road, so it doesn't hurt to prepare

OpenStudy (anonymous):

You're right there... Thanks again!!!

jimthompson5910 (jim_thompson5910):

sure thing